OpenStudy (yourboydusty):

What is the equation of the line containing the points (7, 3), (14, 6), and (21, 9)? A. y = 3/7x B. y = x - 4 C. y =1/7 x i haven't even seen a question like this before.

OpenStudy (yourboydusty):

i am just so confused

OpenStudy (will.h):

Super easy question. All you gotta do is to find the slope of the coordinates they gave you pick 2 points out of the 3 they gave you let's say (14,6) (21,9) and find the slope using the slope formula \[\frac{ y2-y1 }{ x2-x1 }\] substitute the right values in the right positions \[\frac{9-6 }{ 21-14 }\] which would equal 3/7 now see which equation has a slope of 3/7? it is the 1st one y = 3/7x

OpenStudy (will.h):

any questions?

OpenStudy (yourboydusty):

no thank you :)